OpenStudy (anonymous):

(v+7)^2 rewrite without parentheses and simplify

OpenStudy (kc_kennylau):

\[(a+b)^2=a^2+2ab+b^2\]

hero (hero):

(x + 7)^2 = (x + 7)(x + 7) = x(x + 7) + 7(x + 7) = x^2 + 7x + 7x + 49 Finish the rest